OpenStudy (anonymous):

how do you integrate -sin(x)^2

OpenStudy (anonymous):

\[\int\limits_{}^{} -\sin ^{2}(x)\]

OpenStudy (turingtest):

use\[\sin^2x=\frac12(1-\cos(2x))\]

OpenStudy (anonymous):

double angle right?

OpenStudy (anonymous):

once you substitute, you should get \[\int{\frac{cos(2x)}{2}-\frac{1}{2}dx}\] which can be rewritten as \[\int{\frac{cos(2x)}{2}dx-\int\frac{1}{2}dx}\] Since an integral is a linear operator, you can separate it into two different integrals. The second integral is a straightforward one. The first one is a u-substitution. It's usually easier to choose your u to be what is 'inside' the function. Remember that \[\int{cos(x)dx}=-sin(x)+c\]

OpenStudy (anonymous):

alright, thats got it. Thanks